OpenStudy (anonymous):

3/13+(-4/7) perform the indicated operation

jimthompson5910 (jim_thompson5910):

3/13+(-4/7) is the same as 3/13-4/7

jimthompson5910 (jim_thompson5910):

Now find the LCD. What is the LCD?

OpenStudy (anonymous):

(3/13)+(-4/7) (3/13) - (4/7) taking L.C.M=7*13 since they do not have any common multiples. [(7*3)-(13*4)]/(7*13) (21-52)/91 -31/91 = -0.34

OpenStudy (anonymous):

thx so much! i was racking my brain!!